OpenStudy (anonymous):

if f(x)=x and g(x)=1-2x , find the value of f(g(-3))

OpenStudy (anonymous):

g(-3)=1-2(-3)=1+6=7 and hence f(g(-3))=f(7)=7 @bobtopher123

OpenStudy (anonymous):

Thank you @Princer_Jones